The following data relates to Alpha Company.

Units in beginning inventory
Units produced 26,000
Units sold ($300 per unit) 21,000
Variable costs per unit:
Direct materials $35
Direct labor 70
Variable overhead 30
Fixed costs:
Fixed overhead per unit produced $45
Fixed selling and administrative expenses 160,000


Determine the value of ending inventory under variable costing.

a.$550,000

b.$900,000

c.$675,000

d.$525,000
Answer #1

Inventory (Units) = Opening + Units produced - Closing

Inventory = 0+26000-21000 = 5000

Per Unit value of Inventory = Direct Material + Direct wages + Variable Overhead

Per Unit value of Inventory = 35+70+30 = 135

Total Inventory value under variable costing method = 5000*135 = 675000

Answer C is correct